# Logarithmic regret in the ergodic Avellaneda–Stoikov market making model

### v1

@@ A.14. Proof of Proposition 22
By (29), the linear ODE for $t\mapsto v_{r}^{\psi^{\kappa}}(t,q;T;\kappa^{\star})$ can be written in a matrix form. Let $\bm{v}_{r}(t)=[v_{r}^{\psi^{\kappa}}(t,\bar{q};T;\kappa^{\star}),...,v_{r}^{\psi^{\kappa}}(t,\underline{q};T;\kappa^{\star})]$ be an $n-dim$ vector, where $n=\bar{q}-\underline{q}+1$. Now, let $\bm{\tilde{A}}_{r}$ denote an $n$-square matrix whose rows are labelled from $\bar{q}$ to $\underline{q}$ and entries are given by
(47) $\bm{\tilde{A}}_{r}(i,q)=\begin{cases}-(r+\lambda^{+}e^{-\kappa^{\star}\psi^{\kappa,+}(q)}\bm{1}_{q>\underline{q}}+\lambda^{-}e^{-\kappa^{\star}\psi^{\kappa,-}(q)}\bm{1}_{q<\bar{q}}),&\text{if $i=q$},\\ \lambda^{+}e^{-\kappa^{\star}\psi^{\kappa,+}(q)},&\text{if $i=q+1$},\\ \lambda^{-}e^{-\kappa^{\star}\psi^{\kappa,-}(q)},&\text{if $i=q-1$},\\ 0,&\text{otherwise. }\end{cases}$
@@
(48) $b_{i}=\lambda^{+}\psi^{\kappa,+}(i)e^{-\kappa^{\star}\psi^{\kappa,+}(i)}\bm{1}_{i>\underline{q}}+\lambda^{-}\psi^{\kappa,-}(i)e^{-\kappa^{\star}\psi^{\kappa,-}(i)}\bm{1}_{i<\bar{q}}-\phi i^{2},$
@@
where $\bm{W}$ is the $n-$square matrix with only $1$ on the first diagonal element and 00 otherwise. By (30), we know $\gamma(\kappa;\kappa^{\star})\mathds{1}=\bm{U}\bm{W}\bm{U}^{-1}\bm{\tilde{b}}$ with $\mathds{1}$ the $n-$dim vector with all entries $1$. ∎
A.15. Proof of Lemma 24
@@
First, we would like to prove that $\kappa\mapsto\psi^{\kappa}$ with $\psi^{\kappa}$ given by 18 is $C^{\infty}([\underline{K},\bar{K}])$. By Proposition 8 and Theorem 9, it is equivalent to show that $\kappa\mapsto\bm{\hat{\omega}}(\kappa)$ by (19) is $C^{\infty}([\underline{K},\bar{K}])$. Let us consider a matrix $\bm{D}$ as
@@
Proposition 22 gives an analytical solution to $\gamma(\kappa;\kappa^{\star})$ but it is not enough to show $\kappa\mapsto\gamma(\kappa;\kappa^{\star})$ is twice differentiable, as $\bm{U}$’s columns are the eigenvectors of $\bm{\tilde{A}}_{0}$ whereas $\bm{\tilde{A}}_{0}$ is not a self-adjoint or normal operator. Let us consider the matrix $\bm{V}$ as
@@
with $v_{q}=\frac{\prod_{i=\bar{q}-1}^{q}\sqrt{\lambda^{-}}e^{-\frac{1}{2}\kappa^{\star}\psi^{\kappa,-}(i)}}{\prod_{i=\bar{q}}^{q+1}\sqrt{\lambda^{+}}e^{-\frac{1}{2}\kappa^{\star}\psi^{\kappa,+}(i)}}$ for $q\in\{\bar{q}-1,\bar{q}-2,\dots,\underline{q}\}$ and $v_{\bar{q}}=1$. Then $\bm{J}=\bm{V}^{-1}\bm{\tilde{A}}_{0}\bm{V}$ is a real and symmetric tridiagonal matrix with the entries as
(55) $\bm{J}(i,q)=\begin{cases}-(\lambda^{+}e^{-\kappa^{\star}\psi^{\kappa,+}(q)}\bm{1}_{q>\underline{q}}+\lambda^{-}e^{-\kappa^{\star}\psi^{\kappa,-}(q)}\bm{1}_{q<\bar{q}}),&\text{if $i=q$},\\ \sqrt{\lambda^{+}\lambda^{-}}e^{-\frac{\kappa^{\star}}{\kappa}},&\text{if $i=q+1$},\\ \sqrt{\lambda^{+}\lambda^{-}}e^{-\frac{\kappa^{\star}}{\kappa}},&\text{if $i=q-1$},\\ 0,&\text{otherwise. }\end{cases}$
@@
$\gamma(\kappa;\kappa^{\star})\mathds{1}=\bm{V}\bm{U^{\prime}}\bm{W}\bm{U^{\prime}}^{-1}\bm{V}^{-1}\bm{\tilde{b}}.$
@@
(56) $\gamma(\kappa;\kappa^{\star})\mathds{1}=\bm{V}Diag(\bm{u}_{1})\bm{W^{\prime}}Diag(\bm{u}_{1})\bm{V}^{-1}\bm{\tilde{b}}.$
Clearly, $\psi^{\kappa}\mapsto e^{-\kappa^{\star}\psi^{\kappa}}$ is a smooth function and $\kappa\mapsto\psi^{\kappa}$ is $C^{\infty}([\underline{K},\bar{K}])$. Therefore, $\bm{V}(\kappa)$ (54), $\bm{V}^{-1}(\kappa)$, $\bm{\tilde{b}}(\kappa)$ (48) and $\bm{J}(\kappa)$ are $C^{\infty}([\underline{K},\bar{K}])$. Moreover, $\bm{J}$ is a real and symmetric tridiagonal matrix with the simple eigenvalue $\lambda_{1}$. By [5, Result 7.2; Theorem 7.6], we know that $\kappa\mapsto\bm{u}_{1}(\kappa)$ can be chosen to be parameterised smoothly in $\kappa$. Hence, by (56), $\kappa\mapsto\gamma(\kappa;\kappa^{\star})$ is at least $C^{2}([\underline{K},\bar{K}])$ and $\frac{\mathrm{d}^{2}}{\mathrm{d}\kappa^{2}}_{\kappa}\gamma(\kappa;\kappa^{\star})$ is bounded on the compact set $\kappa\in[\underline{K},\bar{K}]$. ∎
@@ A.16. Proof of Proposition 26
$\begin{split}\gamma(\kappa;\kappa^{\star})=-\phi q^{2}+&\lambda^{+}e^{-\kappa^{\star}\psi^{\kappa,+}(q)}\Big{(}\hat{v}^{\psi^{\kappa}}(q-1;\kappa^{\star})-\hat{v}^{\psi^{\kappa}}(q;\kappa^{\star})+\psi^{\kappa,+}(q)\Big{)}\bm{1}_{q>\underline{q}}\\ +&\lambda^{-}e^{-\kappa^{\star}\psi^{\kappa,-}(q)}\Big{(}\hat{v}^{\psi^{\kappa}}(q+1;\kappa^{\star})-\hat{v}^{\psi^{\kappa}}(q;\kappa^{\star})+\psi^{\kappa,-}(q)\Big{)}\bm{1}_{q<\bar{q}}\,.\end{split}$
@@
Take integral from 00 to $T$ and then take the expectation with respect to the probability measure $\pi^{\psi^{\kappa};\kappa^{\star}}$ under the controlled SDE (22) with $Q_{0}\sim\pi^{\psi^{\kappa};\kappa^{\star}}$, we have
$\int_{\Omega^{Q}}\int_{0}^{T}\gamma(\kappa;\kappa^{\star})\,\mathrm{d}t\,\mathrm{d}\pi^{\psi^{\kappa};\kappa^{\star}}=\int_{\Omega^{Q}}\int_{0}^{T}\Big{(}\lambda^{+}\psi^{\kappa,+}(q)e^{-\kappa^{\star}\psi^{\kappa,+}(q)}+\lambda^{-}\psi^{\kappa,-}(q)e^{-\kappa^{\star}\psi^{\kappa,-}(q)}$
$-\phi q^{2}+\lambda^{+}e^{-\kappa^{\star}\psi^{\kappa,+}}\big{(}\hat{v}^{\psi^{\kappa}}(q-1;\kappa^{\star})-\hat{v}^{\psi^{\kappa}}(q;\kappa^{\star})\big{)}\bm{1}_{q>\underline{q}}$
$+\lambda^{-}e^{-\kappa^{\star}\psi^{\kappa,-}}\big{(}\hat{v}^{\psi^{\kappa}}(q+1;\kappa^{\star})-\hat{v}^{\psi^{\kappa}}(q;\kappa^{\star})\big{)}\bm{1}_{q<\bar{q}}\Big{)}\,\mathrm{d}t\,\mathrm{d}\pi^{\psi^{\kappa};\kappa^{\star}}\,,$
@@
where we omit the indicator function in the first line since $\psi^{\kappa,\pm}(q)e^{-\kappa^{\star}\psi^{\kappa,\pm}(q)}=0$ when $q=\bar{q},\underline{q}$, respectively. As $\gamma(\kappa;\kappa^{\star})$ is independent of $q$ and $t$ by Proposition 21, dividing by $T>0$ we get
@@
$\begin{split}&\gamma(\kappa;\kappa^{\star})\\ &=\frac{1}{T}\mathbb{E}_{\pi^{\psi^{\kappa};\kappa^{\star}}}\Big{[}\int_{0}^{T}\lambda^{+}\psi^{\kappa,+}e^{-\kappa^{\star}\psi^{\kappa,+}(Q_{t}^{\psi^{\kappa};\kappa^{\star}})}+\lambda^{-}\psi^{\kappa,-}e^{-\kappa^{\star}\psi^{\kappa,-}(Q_{t}^{\psi^{\kappa};\kappa^{\star}})}-\phi(Q_{t}^{\psi^{\kappa};\kappa^{\star}})^{2}\,\mathrm{d}t\Big{]}\\ &\qquad+\frac{1}{T}\int_{\Omega^{Q}}\int_{0}^{T}\Big{(}\lambda^{+}e^{-\kappa^{\star}\psi^{\kappa,+}(q)}\big{(}\hat{v}^{\psi^{\kappa}}(q-1;\kappa^{\star})-\hat{v}^{\psi^{\kappa}}(q;\kappa^{\star})\big{)}\bm{1}_{q>\underline{q}}\\ &\qquad\qquad+\lambda^{-}e^{-\kappa^{\star}\psi^{\kappa,-}(q)}\big{(}\hat{v}^{\psi^{\kappa}}(q+1;\kappa^{\star})-\hat{v}^{\psi^{\kappa}}(q;\kappa^{\star})\big{)}\bm{1}_{q<\bar{q}}\Big{)}\,\mathrm{d}t\,\mathrm{d}\pi^{\psi^{\kappa};\kappa^{\star}}.\end{split}$
Moreover, for the initial distribution $\pi^{\psi^{\kappa};\kappa^{\star}}$, we have
@@
$\begin{split}&\gamma(\kappa;\kappa^{\star})=\lim_{T\rightarrow+\infty}\frac{1}{T}v_{0}^{\psi^{\kappa}}(t=0,q;T;\kappa^{\star})\\ &=\lim_{T\rightarrow+\infty}\frac{1}{T}\mathbb{E}_{\pi^{\psi^{\kappa};\kappa^{\star}}}\Big{[}\int_{0}^{T}\Big{(}\lambda^{+}\psi^{\kappa,+}e^{-\kappa^{\star}\psi^{\kappa,+}}+\lambda^{-}\psi^{\kappa,-}e^{-\kappa^{\star}\psi^{\kappa,-}}-\phi(Q^{\psi^{\kappa};\kappa^{\star}}_{t})^{2}\Big{)}\,\mathrm{d}t\Big{]}\end{split}$
@@
$\begin{split}0=\int_{\Omega^{Q}}\lambda^{+}e^{-\kappa^{\star}\psi^{\kappa,+}}\Big{(}&\hat{v}^{\psi^{\kappa}}(q-1;\kappa^{\star})-\hat{v}^{\psi^{\kappa}}(q;\kappa^{\star})\Big{)}\bm{1}_{q>\underline{q}}\\ &+\lambda^{-}e^{-\kappa^{\star}\psi^{\kappa,-}}\Big{(}\hat{v}^{\psi^{\kappa}}(q+1;\kappa^{\star})-\hat{v}^{\psi^{\kappa}}(q;\kappa^{\star})\Big{)}\bm{1}_{q<\bar{q}}\,\mathrm{d}\pi^{\psi^{\kappa};\kappa^{\star}},\end{split}$

### v2

@@ A.13. Proof of Proposition 20
By (33), the linear ODE for $t\mapsto v_{r}^{\psi^{\kappa}}(t,q;T;\kappa^{\ast})$ can be written in a matrix form. Let $\bm{v}_{r}(t)=[v_{r}^{\psi^{\kappa}}(t,\bar{q};T;\kappa^{\ast}),...,v_{r}^{\psi^{\kappa}}(t,\underline{q};T;\kappa^{\ast})]$ be an $n-dim$ vector, where $n=\bar{q}-\underline{q}+1$. Now, let $\bm{\tilde{A}}_{r}$ denote an $n$-square matrix whose rows are labelled from $\bar{q}$ to $\underline{q}$ and entries are given by
(50) $\bm{\tilde{A}}_{r}(i,q)=\begin{cases}-(r+\lambda^{+}e^{-\kappa^{\ast}\psi^{\kappa,+}(q)}\mathds{1}_{q>\underline{q}}+\lambda^{-}e^{-\kappa^{\ast}\psi^{\kappa,-}(q)}\mathds{1}_{q<\bar{q}}),&\text{if $i=q$},\\ \lambda^{+}e^{-\kappa^{\ast}\psi^{\kappa,+}(q)},&\text{if $i=q+1$},\\ \lambda^{-}e^{-\kappa^{\ast}\psi^{\kappa,-}(q)},&\text{if $i=q-1$},\\ 0,&\text{otherwise. }\end{cases}$
@@
(51) $b_{i}=\lambda^{+}\psi^{\kappa,+}(i)e^{-\kappa^{\ast}\psi^{\kappa,+}(i)}\mathds{1}_{i>\underline{q}}+\lambda^{-}\psi^{\kappa,-}(i)e^{-\kappa^{\ast}\psi^{\kappa,-}(i)}\mathds{1}_{i<\bar{q}}-\phi i^{2},$
@@
where $\bm{W}$ is the $n-$square matrix with only $1$ on the first diagonal element and 00 otherwise. By (34), we know $\gamma(\kappa;\kappa^{\ast})\mathds{1}=\bm{U}\bm{W}\bm{U}^{-1}\bm{\tilde{b}}$ with $\mathds{1}$ the $n-$dim vector with all entries $1$. ∎
A.14. Proof of Lemma 23
@@
First, we would like to prove that $\kappa\mapsto\psi^{\kappa}$ with $\psi^{\kappa}$ given by (19) is $C^{\infty}([\underline{K},\bar{K}])$. By Proposition 8 and Theorem 9, it is equivalent to show that $\kappa\mapsto\bm{\hat{\omega}}(\kappa)$ by (20) is $C^{\infty}([\underline{K},\bar{K}])$. Let us consider a matrix $\bm{D}$ as
@@
Proposition 20 gives an analytical solution to $\gamma(\kappa;\kappa^{\ast})$ but it is not enough to show $\kappa\mapsto\gamma(\kappa;\kappa^{\ast})$ is twice differentiable, as $\bm{U}$’s columns are the eigenvectors of $\bm{\tilde{A}}_{0}$ whereas $\bm{\tilde{A}}_{0}$ is not a self-adjoint or normal operator. Let us consider the matrix $\bm{V}$ as
@@
with $v_{q}=\frac{\prod_{i=\bar{q}-1}^{q}\sqrt{\lambda^{-}}e^{-\frac{1}{2}\kappa^{\ast}\psi^{\kappa,-}(i)}}{\prod_{i=\bar{q}}^{q+1}\sqrt{\lambda^{+}}e^{-\frac{1}{2}\kappa^{\ast}\psi^{\kappa,+}(i)}}$ for $q\in\{\bar{q}-1,\bar{q}-2,\dots,\underline{q}\}$ and $v_{\bar{q}}=1$. Then $\bm{J}=\bm{V}^{-1}\bm{\tilde{A}}_{0}\bm{V}$ is a real and symmetric tridiagonal matrix with the entries as
(58) $\bm{J}(i,q)=\begin{cases}-(\lambda^{+}e^{-\kappa^{\ast}\psi^{\kappa,+}(q)}\mathds{1}_{q>\underline{q}}+\lambda^{-}e^{-\kappa^{\ast}\psi^{\kappa,-}(q)}\mathds{1}_{q<\bar{q}}),&\text{if $i=q$},\\ \sqrt{\lambda^{+}\lambda^{-}}e^{-\frac{\kappa^{\ast}}{\kappa}},&\text{if $i=q+1$},\\ \sqrt{\lambda^{+}\lambda^{-}}e^{-\frac{\kappa^{\ast}}{\kappa}},&\text{if $i=q-1$},\\ 0,&\text{otherwise. }\end{cases}$
@@
$\gamma(\kappa;\kappa^{\ast})\mathds{1}=\bm{V}\bm{U^{\prime}}\bm{W}\bm{U^{\prime}}^{-1}\bm{V}^{-1}\bm{\tilde{b}}.$
@@
(59) $\gamma(\kappa;\kappa^{\ast})\mathds{1}=\bm{V}Diag(\bm{u}_{1})\bm{W^{\prime}}Diag(\bm{u}_{1})\bm{V}^{-1}\bm{\tilde{b}}.$
Clearly, $\psi^{\kappa}\mapsto e^{-\kappa^{\ast}\psi^{\kappa}}$ is a smooth function and $\kappa\mapsto\psi^{\kappa}$ is $C^{\infty}([\underline{K},\bar{K}])$. Therefore, $\bm{V}(\kappa)$ (57), $\bm{V}^{-1}(\kappa)$, $\bm{\tilde{b}}(\kappa)$ (51) and $\bm{J}(\kappa)$ are $C^{\infty}([\underline{K},\bar{K}])$. Moreover, $\bm{J}$ is a real and symmetric tridiagonal matrix with the simple eigenvalue $\lambda_{1}$. By [5, Result 7.2; Theorem 7.6], we know that $\kappa\mapsto\bm{u}_{1}(\kappa)$ can be chosen to be parameterised smoothly in $\kappa$. Hence, by (59), $\kappa\mapsto\gamma(\kappa;\kappa^{\ast})$ is at least $C^{2}([\underline{K},\bar{K}])$ and $\frac{\mathrm{d}^{2}}{\mathrm{d}\kappa^{2}}\gamma(\kappa;\kappa^{\ast})$ is bounded on the compact set $\kappa\in[\underline{K},\bar{K}]$. ∎
A.15. Proof of Lemma 27
Proof.
The state process $\big{(}Q_{t}^{\psi^{\kappa};\kappa^{\ast}}\big{)}_{t\geq 0}$ (23) is driven by two independent Poisson jump processes with intensities $\lambda^{+}e^{-\kappa^{\ast}\delta^{+}}$ and $\lambda^{-}e^{-\kappa^{\ast}\delta^{-}}$, respectively. The depths $\delta^{\pm}$ are uniquely and continuously determined by the ergodic optimal control $q\mapsto\psi^{\kappa}(q)$ given the agent’s current position $q=Q_{t}$ at time $t$, where $\psi^{\kappa}(q)$ given by (19) is the ergodic optimal control under the misspecified parameter $\kappa$. As a result, the transition probability at any time $t>0$ depends only on the current state, implying that the stochastic process $(Q_{t})_{t\geq 0}$–where we omit the superscript for notational simplicity–satisfies the Markov property. Furthermore, the state space $\Omega^{Q}=[\underline{q},\bar{q}]\cap\mathbb{Z}$ is discrete and finite. Hence $(Q_{t})_{t\geq 0}$ can be equivalently represented as a continuous-time Markov chain with a finite state space.
Let $\bm{Q}=(\bm{Q}_{ij})_{i,j\in\Omega^{Q}}$ denote the transition rate matrix, where the indices are labelled from $\bar{q}$ to $\underline{q}$. Each entry $\bm{Q}_{ij}$ represents the instantaneous transition rate of the process from state $i$ to state $j$, which can be derived from the infinitesimal generator of $(Q_{t})_{t\geq 0}$. Hence the entries of $\bm{Q}$ are
(60) $\bm{Q}_{ij}=\begin{cases}-\big{(}\lambda^{+}e^{-\kappa^{\ast}\psi^{\kappa,+}(i)}\mathds{1}_{i>\underline{q}}+\lambda^{-}e^{-\kappa^{\ast}\psi^{\kappa,-}(i)}\mathds{1}_{i<\bar{q}}\big{)},&\text{if $i=j$},\\ \lambda^{-}e^{-\kappa^{\ast}\psi^{\kappa,-}(i)},&\text{if $i=j-1$},\\ \lambda^{+}e^{-\kappa^{\ast}\psi^{\kappa,+}(i)},&\text{if $i=j+1$},\\ 0,&\text{otherwise}.\end{cases}$
From (60), one may notice that $\big{(}Q_{t}^{\psi^{\kappa};\kappa^{\ast}}\big{)}_{t\geq 0}$ is equivalent to a general Birth-Death process with a finite state space. Hence there exists a unique equilibrium distribution $\pi$, for $\big{(}Q_{t}^{\psi^{\kappa};\kappa^{\ast}}\big{)}_{t\geq 0}$ when $t$ goes to infinity [37, Theorem 5.5.3]. Moreover, $\pi$ is uniquely determined by
$\pi\bm{Q}=0,\quad\text{subject to }\sum_{q\in\Omega^{Q}}\pi_{q}=1\,.$
@@ A.16. Proof of Proposition 29
$\begin{split}\gamma(\kappa;\kappa^{\ast})=-\phi q^{2}+&\lambda^{+}e^{-\kappa^{\ast}\psi^{\kappa,+}(q)}\Big{(}\hat{v}^{\psi^{\kappa}}(q-1;\kappa^{\ast})-\hat{v}^{\psi^{\kappa}}(q;\kappa^{\ast})+\psi^{\kappa,+}(q)\Big{)}\mathds{1}_{q>\underline{q}}\\ +&\lambda^{-}e^{-\kappa^{\ast}\psi^{\kappa,-}(q)}\Big{(}\hat{v}^{\psi^{\kappa}}(q+1;\kappa^{\ast})-\hat{v}^{\psi^{\kappa}}(q;\kappa^{\ast})+\psi^{\kappa,-}(q)\Big{)}\mathds{1}_{q<\bar{q}}\,.\end{split}$
@@
Take integral from 00 to $T$ and then take the expectation with respect to the probability measure $\pi^{\psi^{\kappa};\kappa^{\ast}}$ under the controlled SDE (23) with $Q_{0}\sim\pi^{\psi^{\kappa};\kappa^{\ast}}$, we have
$\int_{\Omega^{Q}}\int_{0}^{T}\gamma(\kappa;\kappa^{\ast})\,\mathrm{d}t\,\mathrm{d}\pi^{\psi^{\kappa};\kappa^{\ast}}=\int_{\Omega^{Q}}\int_{0}^{T}\Big{(}\lambda^{+}\psi^{\kappa,+}(q)e^{-\kappa^{\ast}\psi^{\kappa,+}(q)}+\lambda^{-}\psi^{\kappa,-}(q)e^{-\kappa^{\ast}\psi^{\kappa,-}(q)}$
$-\phi q^{2}+\lambda^{+}e^{-\kappa^{\ast}\psi^{\kappa,+}}\big{(}\hat{v}^{\psi^{\kappa}}(q-1;\kappa^{\ast})-\hat{v}^{\psi^{\kappa}}(q;\kappa^{\ast})\big{)}\mathds{1}_{q>\underline{q}}$
$+\lambda^{-}e^{-\kappa^{\ast}\psi^{\kappa,-}}\big{(}\hat{v}^{\psi^{\kappa}}(q+1;\kappa^{\ast})-\hat{v}^{\psi^{\kappa}}(q;\kappa^{\ast})\big{)}\mathds{1}_{q<\bar{q}}\Big{)}\,\mathrm{d}t\,\mathrm{d}\pi^{\psi^{\kappa};\kappa^{\ast}}\,,$
@@
where we omit the indicator function in the first line since $\psi^{\kappa,\pm}(q)e^{-\kappa^{\ast}\psi^{\kappa,\pm}(q)}=0$ when $q=\bar{q},\underline{q}$, respectively. As $\gamma(\kappa;\kappa^{\ast})$ is independent of $q$ and $t$ by Proposition 19, dividing by $T>0$ we get
@@
$\begin{split}&\gamma(\kappa;\kappa^{\ast})\\ &=\frac{1}{T}\mathbb{E}_{\pi^{\psi^{\kappa};\kappa^{\ast}}}\Big{[}\int_{0}^{T}\lambda^{+}\psi^{\kappa,+}e^{-\kappa^{\ast}\psi^{\kappa,+}(Q_{t}^{\psi^{\kappa};\kappa^{\ast}})}+\lambda^{-}\psi^{\kappa,-}e^{-\kappa^{\ast}\psi^{\kappa,-}(Q_{t}^{\psi^{\kappa};\kappa^{\ast}})}-\phi(Q_{t}^{\psi^{\kappa};\kappa^{\ast}})^{2}\,\mathrm{d}t\Big{]}\\ &\qquad+\frac{1}{T}\int_{\Omega^{Q}}\int_{0}^{T}\Big{(}\lambda^{+}e^{-\kappa^{\ast}\psi^{\kappa,+}(q)}\big{(}\hat{v}^{\psi^{\kappa}}(q-1;\kappa^{\ast})-\hat{v}^{\psi^{\kappa}}(q;\kappa^{\ast})\big{)}\mathds{1}_{q>\underline{q}}\\ &\qquad\qquad+\lambda^{-}e^{-\kappa^{\ast}\psi^{\kappa,-}(q)}\big{(}\hat{v}^{\psi^{\kappa}}(q+1;\kappa^{\ast})-\hat{v}^{\psi^{\kappa}}(q;\kappa^{\ast})\big{)}\mathds{1}_{q<\bar{q}}\Big{)}\,\mathrm{d}t\,\mathrm{d}\pi^{\psi^{\kappa};\kappa^{\ast}}.\end{split}$
Moreover, for the initial distribution $\pi^{\psi^{\kappa};\kappa^{\ast}}$, we have
@@
$\begin{split}&\gamma(\kappa;\kappa^{\ast})=\lim_{T\rightarrow+\infty}\frac{1}{T}v_{0}^{\psi^{\kappa}}(0,q;T;\kappa^{\ast})\\ &=\lim_{T\rightarrow+\infty}\frac{1}{T}\mathbb{E}_{\pi^{\psi^{\kappa};\kappa^{\ast}}}\Big{[}\int_{0}^{T}\Big{(}\lambda^{+}\psi^{\kappa,+}e^{-\kappa^{\ast}\psi^{\kappa,+}}+\lambda^{-}\psi^{\kappa,-}e^{-\kappa^{\ast}\psi^{\kappa,-}}-\phi(Q^{\psi^{\kappa};\kappa^{\ast}}_{t})^{2}\Big{)}\,\mathrm{d}t\Big{]}\end{split}$
@@
$\begin{split}0=\int_{\Omega^{Q}}\lambda^{+}e^{-\kappa^{\ast}\psi^{\kappa,+}}\Big{(}&\hat{v}^{\psi^{\kappa}}(q-1;\kappa^{\ast})-\hat{v}^{\psi^{\kappa}}(q;\kappa^{\ast})\Big{)}\mathds{1}_{q>\underline{q}}\\ &+\lambda^{-}e^{-\kappa^{\ast}\psi^{\kappa,-}}\Big{(}\hat{v}^{\psi^{\kappa}}(q+1;\kappa^{\ast})-\hat{v}^{\psi^{\kappa}}(q;\kappa^{\ast})\Big{)}\mathds{1}_{q<\bar{q}}\,\mathrm{d}\pi^{\psi^{\kappa};\kappa^{\ast}},\end{split}$